# Vertex-based Graph Neural Solver and its Application to Linear Elasticity Equations

### v1

@@ 2.2 Graph Representation of Linear Systems
To leverage GNN for solving linear systems, the discrete system (4) must be mapped onto a graph structure $\mathcal{G}=(\mathcal{V},\mathcal{E})$. For discrete systems arising from vector-valued PDEs, we adopt a block-level representation: the computational mesh $\mathcal{T}_{h}$ is treated as the graph, containing $N$ nodes ($|\mathcal{V}|=N$), with adjacency relationships determined by the geometric connectivity of the mesh. This approach naturally preserves the topological structure of the physical space and the coupling information between vector components.
@@
In this representation, the block structure information of the system is encoded into node features $V$ and edge features $E$
@@
  • Node Features: Composed of physical quantities defined on the nodes. We utilize the block structure of the right-hand side term $\boldsymbol{f}$ as node features
    $V=\{\boldsymbol{f}_{i}\in\mathbb{R}^{d}\mid i=1,\cdots,N\}.$ (5)
@@
  • Edge Features: Composed of the non-zero blocks $\boldsymbol{A}_{ij}\in\mathbb{R}^{d\times d}$ of the stiffness matrix. To facilitate GNN processing, these matrix blocks must be flattened into vectors. In this work, we adopt a row-major flattening convention, denoted as $\text{Flatten}(\cdot)$
    $E=\left\{\,\boldsymbol{e}_{ij}=\text{Flatten}(\boldsymbol{A}_{ij})\in\mathbb{R}^{d^{2}}\ :\ (i,j)\in\mathcal{E}\right\}.$ (6)
Block Matrix-Vector Multiplication on Graphs
@@
Within the graph-based framework, sparse Block Matrix-Vector multiplication (Block SpMV) serves as a core operation, exemplified by the block weighted Jacobi iteration
$\boldsymbol{u}^{k+1}=\boldsymbol{u}^{k}+\omega\boldsymbol{D}^{-1}(\boldsymbol{f}-\boldsymbol{A}\boldsymbol{u}^{k}),$ (7)
@@
where $\omega$ is the relaxation factor and $\boldsymbol{D}$ is the block-diagonal part of $\boldsymbol{A}$. The Block SpMV operation $(\boldsymbol{A}\boldsymbol{u})_{i}=\sum_{j}\boldsymbol{A}_{ij}\boldsymbol{u}_{j}$ can be efficiently implemented via a message-passing mechanism on the graph. The specific steps are as follows (using the calculation of $\boldsymbol{w}=\boldsymbol{A}\boldsymbol{v}$ as an example):
@@
    Feature Expansion (Vector Replication): To perform element-wise operations with the row-major flattened edge features $\boldsymbol{e}_{ij}$, the feature vector $\boldsymbol{v}_{j}\in\mathbb{R}^{d}$ of the source node $j$ must be replicated $d$ times and concatenated to obtain the extended feature $\boldsymbol{v}^{\prime}_{j}\in\mathbb{R}^{d^{2}}$:
    $\boldsymbol{v}^{\prime}_{j}=\text{Concat}(\underbrace{\boldsymbol{v}_{j},\boldsymbol{v}_{j},\cdots,\boldsymbol{v}_{j}}_{d\text{ times}}).$ (8)
@@
    Message Calculation (Element-wise Multiplication): For each edge $(i,j)$, the message $\boldsymbol{m}_{ij}$ from source node $j$ to target node $i$ is calculated. This is achieved via the Hadamard product ($\odot$) of the edge feature $\boldsymbol{e}_{ij}$ and the expanded node feature $\boldsymbol{v}^{\prime}_{j}$:
    $\boldsymbol{m}_{ij}=\boldsymbol{e}_{ij}\odot\boldsymbol{v}^{\prime}_{j}\in\mathbb{R}^{d^{2}}.$ (9)
    This operation computes all element-wise product terms required for the matrix multiplication $\boldsymbol{A}_{ij}\boldsymbol{v}_{j}$ in parallel.
@@
    Message Aggregation (Summation): For each target node $i$, messages from all neighboring nodes $j\in\mathcal{N}(i)$ are aggregated:
    $\boldsymbol{h}_{i}=\sum_{j\in\mathcal{N}(i)}\boldsymbol{m}_{ij}\in\mathbb{R}^{d^{2}}.$ (10)
@@
    Result Reconstruction (Intra-block Summation): The aggregated result $\boldsymbol{h}_{i}$ is a $d^{2}$-dimensional vector. To obtain the final $d$-dimensional result $\boldsymbol{w}_{i}$, $\boldsymbol{h}_{i}$ is treated as $d$ blocks of length $d$, and the elements within each block are summed to complete the dot product operation:
    $(\boldsymbol{w}_{i})_{k}=\sum_{p=1}^{d}(\boldsymbol{h}_{i})_{(k-1)d+p},\quad k=1,\dots,d.$ (11)
Through this process, block-structured linear algebraic operations are fully mapped to feature manipulations and message passing on the graph, laying the foundation for constructing GNN-based solvers.
@@ 2.3 Hybrid Iterative Methods and Fourier Neural Solver
Traditional simple iterative methods (such as the block weighted Jacobi method) are typically efficient at eliminating high-frequency errors (the smoothing effect) but are inefficient at reducing low-frequency errors. Conversely, deep neural networks possess a “spectral bias” [39, 40], making them more adept at capturing and correcting global low-frequency information. Based on this complementarity, a Deep Learning-based Hybrid Iterative Method (DL-HIM) [30] can be constructed as follows:
Smoothing: $\boldsymbol{u}^{(k+\frac{1}{2})}=\boldsymbol{u}^{(k)}+\mathcal{B}\bigl(\boldsymbol{f}-\boldsymbol{A}\boldsymbol{u}^{(k)}\bigr),$ (12a)
Correction: $\boldsymbol{u}^{(k+1)}=\boldsymbol{u}^{(k+\frac{1}{2})}+\mathcal{H}\bigl(\boldsymbol{f}-\boldsymbol{A}\boldsymbol{u}^{(k+\frac{1}{2})}\bigr),$ (12b)
where $\mathcal{B}$ is a traditional smoother, and $\mathcal{H}$ is a global corrector based on neural networks.
The Fourier Neural Solver (FNS) [29] proposes a specific structure for $\mathcal{H}$ within this framework. Its design is inspired by the eigendecomposition of the matrix $\boldsymbol{A}=\boldsymbol{Q}\Lambda\boldsymbol{Q}^{-1}$, implying the inverse is $\boldsymbol{A}^{-1}=\boldsymbol{Q}\Lambda^{-1}\boldsymbol{Q}^{-1}$. In specific cases (e.g., constant coefficient PDEs on structured grids), the eigenvector matrix $\boldsymbol{Q}$ can be approximated by the Fourier basis. Based on this observation, FNS constructs the corrector (approximating $\boldsymbol{A}^{-1}$) as:
@@
here, $\mathcal{F}$ denotes the Fast Fourier Transform (from physical space to the frequency domain), and $\mathcal{F}^{-1}$ denotes the inverse transform. $\widetilde{\Lambda}$ is a learnable diagonal matrix used to approximate the inverse of the eigenvalue matrix $\Lambda^{-1}$.
However, for general variable-coefficient problems, there is a significant deviation between the true eigenbasis and the Fourier basis. To overcome this limitation, [41] further improved this method, which introduces a transition matrix $\boldsymbol{T}$ to describe the relationship between these bases, assuming the true eigenbasis can be represented as a transformation of the Fourier basis: $\boldsymbol{Q}=\mathcal{F}^{-1}\boldsymbol{T}$. The improved operator takes the form:
$\mathcal{H}=\boldsymbol{Q}\Lambda^{-1}\boldsymbol{Q}^{*}\approx\mathcal{F}^{-1}\,\mathcal{C}\,\widetilde{\Lambda}\,\mathcal{C}^{*}\,\mathcal{F}$, (14)
where $\mathcal{C}$ is typically designed as a Convolutional Neural Network (CNN) structure to approximate $\boldsymbol{T}$, and $\mathcal{C}^{*}$ represents its conjugate transpose (often corresponding to transposed convolution in CNN implementations).
Although FNS significantly improves performance, its application remains largely limited to structured grids and scalar PDEs. This limitation stems from its reliance on regular geometry and fixed discretization stencils, which ensure that the operator can be represented in convolutional form. Consequently, this work aims to construct a more general FNS that does not depend on specific stencils, is applicable to arbitrary unstructured grids, and can effectively handle vector-valued PDEs (such as linear elasticity problems).

### v2

@@ 2.2 Graph Representation and Message Passing Mechanism
To facilitate the application of GNNs, the discrete system (5) is mapped onto a graph structure $\mathcal{G}=(\mathcal{V},\mathcal{E})$. We employ a block-level representation wherein the computational mesh $\mathcal{T}_{h}$ defines the graph topology ($|\mathcal{V}|=N$), with adjacency determined by the geometric connectivity of the mesh. This approach naturally preserves both the topological structure of the physical domain and the vector coupling inherent in the elasticity equations.
@@
The block structure of the linear system is encoded into node features $V$ and edge features $E$ as follows:
@@
  • Node Features: These consist of physical quantities defined at the nodes. Specifically, we assign the block components of the right-hand side vector $\bm{f}$ as node features:
    $V=\{\bm{f}_{i}\in\mathbb{R}^{d}\mid i=1,\dots,N\}.$ (6)
@@
  • Edge Features: These comprise the non-zero blocks $\bm{A}_{ij}\in\mathbb{R}^{d\times d}$ of the stiffness matrix. To align with standard GNN operations, the matrix blocks are flattened into vectors using a row-major convention:
    $E=\left\{\bm{e}_{ij}=\text{Flatten}(\bm{A}_{ij})\in\mathbb{R}^{d^{2}}\mid(i,j)\in\mathcal{E}\right\}.$ (7)
Graph-Based Block Matrix-Vector Multiplication
@@
Sparse Block Matrix-Vector Multiplication (Block SpMV) constitutes a kernel operation within the graph-based framework, essential for iterative schemes such as the weighted Block Jacobi method:
$\bm{u}^{(k+1)}=\bm{u}^{(k)}+\omega\bm{D}^{-1}(\bm{f}-\bm{A}\bm{u}^{(k)}),$ (8)
@@
where $\omega$ is a relaxation factor and $\bm{D}$ represents the block-diagonal component of $\bm{A}$. The local operation $(\bm{A}\bm{u})_{i}=\sum_{j}\bm{A}_{ij}\bm{u}_{j}$ is implemented via a message-passing mechanism. Taking the computation of $\bm{w}=\bm{A}\bm{v}$ as an example, the procedure involves four steps:
@@
    Feature Expansion: To enable element-wise operations with the flattened edge features $\bm{e}_{ij}$, the feature vector $\bm{v}_{j}\in\mathbb{R}^{d}$ of the source node $j$ is replicated $d$ times and concatenated to form an expanded vector $\bm{v}^{\prime}_{j}\in\mathbb{R}^{d^{2}}$:
    $\bm{v}^{\prime}_{j}=\text{Concat}(\underbrace{\bm{v}_{j},\bm{v}_{j},\dots,\bm{v}_{j}}_{d\text{ times}}).$ (9)
@@
    Message Computation: For each edge $(i,j)$, the message $\bm{m}_{ij}$ is computed via the Hadamard product ($\odot$) of the edge feature $\bm{e}_{ij}$ and the expanded source feature $\bm{v}^{\prime}_{j}$:
    $\bm{m}_{ij}=\bm{e}_{ij}\odot\bm{v}^{\prime}_{j}\in\mathbb{R}^{d^{2}}.$ (10)
    This step effectively calculates all component-wise products required for the matrix-vector multiplication $\bm{A}_{ij}\bm{v}_{j}$ in parallel.
@@
    Message Aggregation: Messages from all neighboring nodes $j\in\mathcal{N}(i)$ are aggregated at the target node $i$ via summation:
    $\bm{h}_{i}=\sum_{j\in\mathcal{N}(i)}\bm{m}_{ij}\in\mathbb{R}^{d^{2}}.$ (11)
@@
    Result Reconstruction: The aggregated vector $\bm{h}_{i}$ is reduced to the final $d$-dimensional result $\bm{w}_{i}$ by summing specific components. Viewing $\bm{h}_{i}$ as a concatenation of $d$ segments of length $d$, the $k$-th component of $\bm{w}_{i}$ is obtained by:
    $(\bm{w}_{i})_{k}=\sum_{p=1}^{d}(\bm{h}_{i})_{(k-1)d+p},\quad k=1,\dots,d.$ (12)
This procedure effectively maps block-structured linear algebraic operations onto the GNN feature space, establishing a foundation for the proposed learning-based solver.
3 Deep Learning-Based Hybrid Iterative Methods
This section establishes the theoretical motivation and analytical framework for the proposed method. We rigorously quantify the limitations of classical smoothers in capturing the spectral properties of vector-valued elasticity equations via local Fourier analysis (LFA) [brandt1977multi], thereby justifying the necessity of a global correction operator. Subsequently, we introduce preconditioners based on neural operators and address the challenges imposed by spectral bias. Finally, we formulate the hybrid iterative strategy and provide a rigorous analysis of its convergence properties.
3.1 Local Fourier Analysis of Weighted Block Jacobi Method
We consider the 2D linear elasticity equations for a homogeneous, isotropic medium. Discretizing the system on an infinite uniform grid with mesh size $h$ yields the stencil operator $\mathcal{L}_{h}$. The error is expanded in Fourier modes $\bm{\varphi}(\bm{\theta},\bm{x})=e^{i\bm{\theta}\cdot\bm{x}/h}$, where $\bm{\theta}=(\theta_{1},\theta_{2})\in[-\pi,\pi)^{2}$ denotes the frequency vector.
Applying the discrete operator to these modes yields the symbol matrix $\widetilde{\mathcal{L}}_{h}(\bm{\theta})\in\mathbb{C}^{2\times 2}$, which characterizes the operator’s action in the frequency domain. For standard bilinear finite elements or the equivalent nine-point difference scheme, the symbol matrix explicitly takes the form:
$\widetilde{\mathcal{L}}_{h}(\bm{\theta})=\frac{4}{h^{2}}\begin{bmatrix}(\lambda+2\mu)s_{1}^{2}+\mu s_{2}^{2}&(\lambda+\mu)s_{1}s_{2}c_{1}c_{2}\\[6.0pt] (\lambda+\mu)s_{1}s_{2}c_{1}c_{2}&\mu s_{1}^{2}+(\lambda+2\mu)s_{2}^{2}\end{bmatrix}$, (13)
where we employ the shorthand notation $s_{k}=\sin(\theta_{k}/2)$ and $c_{k}=\cos(\theta_{k}/2)$ for $k=1,2$.
The spectral properties of $\widetilde{\mathcal{L}}_{h}(\bm{\theta})$ are critical for convergence analysis. The matrix is symmetric and positive definite, possessing two eigenvalues $\eta_{\text{vol}}(\bm{\theta})$ and $\eta_{\text{shear}}(\bm{\theta})$, which correspond to the volumetric (curl-free) and shear (divergence-free) modes, respectively. Asymptotically, for low frequencies ($|\bm{\theta}|\to 0$), these eigenvalues scale as:
$\eta_{\text{vol}}(\bm{\theta})\approx\frac{|\bm{\theta}|^{2}}{h^{2}}(\lambda+2\mu),\qquad\eta_{\text{shear}}(\bm{\theta})\approx\frac{|\bm{\theta}|^{2}}{h^{2}}\mu.$ (14)
The symbol of the error propagation operator for the weighted block Jacobi smoother is given by $\widetilde{\mathcal{M}}(\bm{\theta})=\bm{I}-\omega\widetilde{\mathcal{D}}^{-1}\widetilde{\mathcal{L}}_{h}(\bm{\theta})$, where $\widetilde{\mathcal{D}}$ represents the symbol of the block diagonal preconditioner. Crucially, the diagonal entries of the stiffness matrix are dominated by the largest material parameter, scaling as $\widetilde{\mathcal{D}}\approx C(\lambda+2\mu)\bm{I}$.
This formulation reveals the fundamental pathology of classical smoothers in near-incompressible regimes (where $\nu\to 0.5$, implying $\lambda\gg\mu$). While the smoothing factor for the volumetric mode remains well-bounded, the smoothing factor for the shear mode behaves as:
$\rho_{\text{shear}}(\bm{\theta})\approx\left|1-\omega\frac{\eta_{\text{shear}}}{\eta_{\text{diag}}}\right|\approx\left|1-\omega\frac{\mu}{\lambda+2\mu}\right|.$ (15)
As $\nu\to 0.5$, the ratio $\frac{\mu}{\lambda+2\mu}\to 0$, causing $\rho_{\text{shear}}\to 1$. Consequently, high-frequency shear errors undergo minimal attenuation. This spectral disparity—where the smoother is effectively tuned for the stiffest mode (volumetric) but fails to dampen the flexible mode (shear)—explains the stagnation observed in Figure 1.
Figure 1: Spectral radius distributions obtained via LFA with (a) $\nu=0.3$; (b) $\nu=0.4$; (c) $\nu=0.45$.
This analysis confirms that, unlike the scalar Poisson equation, elasticity problems suffer from inherent stiffness disparity between physical modes. This challenge necessitates adaptive correction mechanisms capable of addressing these distinct error components beyond simple low-frequency elimination.
3.2 Neural Operator Preconditioning
The theoretical justification for learning inverse operators rests on the well-posedness of the continuous problem. For vector-valued elasticity equations, Korn’s inequality serves as the fundamental estimate, analogous to the Poincaré inequality in scalar PDEs.
Theorem 1 (Korn’s inequality [ciarlet1988mathematical]).
There exists a constant $C_{K}>0$, depending only on the domain $\Omega$, such that for all $\bm{v}\in[H^{1}(\Omega)]^{d}$ satisfying appropriate boundary conditions
$\int_{\Omega}|\bm{\varepsilon}(\bm{v})|^{2}\,d\bm{x}+\int_{\Omega}|\bm{v}|^{2}\,d\bm{x}\geq C_{K}\,\|\bm{v}\|_{H^{1}}^{2}.$ (16)
This inequality guarantees the equivalence between the energy norm $\|\bm{v}\|_{a}:=\sqrt{a(\bm{v},\bm{v})}$ induced by the bilinear form (4) and the Sobolev norm $\|\bm{v}\|_{H^{1}}$. By the Lax–Milgram theorem, the operator $\mathcal{A}:\bm{V}\to\bm{V}^{\prime}$ is an isomorphism with uniform coercivity and continuity. Theoretically, this ensures that the inverse operator $\mathcal{A}^{-1}$ is well-defined and approximable.
In the context of continuous operator preconditioning, we seek to construct a parameterized neural operator $\mathcal{M}_{\theta}:\bm{V}^{\prime}\to\bm{V}$ approximating $\mathcal{A}^{-1}$. We proceed with the following assumption regarding approximation capability.
Assumption 3.1 (Approximation capability).
Based on the universal approximation theorem for neural operators [chen1995universal, lu2021learning], for any $\varepsilon>0$, there exists a parameter set $\theta$ such that
$\|\mathcal{M}_{\theta}-\mathcal{A}^{-1}\|_{\bm{V}^{\prime}\to\bm{V}}\leq\varepsilon.$ (17)
Theorem 2.
Suppose Assumption 3.1 holds and the bilinear form $a(\cdot,\cdot)$ is uniformly coercive and continuous (guaranteed by Korn’s inequality). If $\varepsilon$ is sufficiently small, there exist mesh-independent constants $c_{1},c_{2}>0$ such that for any $\bm{f}\in\bm{V}^{\prime}$
$c_{1}\,\langle\mathcal{A}^{-1}\bm{f},\bm{f}\rangle\leq\langle\mathcal{M}_{\theta}\bm{f},\bm{f}\rangle\leq c_{2}\,\langle\mathcal{A}^{-1}\bm{f},\bm{f}\rangle.$ (18)
Proof.
Decompose the neural operator as $\mathcal{M}_{\theta}=\mathcal{A}^{-1}+\mathcal{E}$, where $\|\mathcal{E}\|\leq\varepsilon$. The quadratic form is given by
$\langle\mathcal{M}_{\theta}\bm{f},\bm{f}\rangle=\langle\mathcal{A}^{-1}\bm{f},\bm{f}\rangle+\langle\mathcal{E}\bm{f},\bm{f}\rangle.$ (19)
Using the coercivity of $\mathcal{A}$ and the boundedness of $\mathcal{E}$, we estimate
$|\langle\mathcal{E}\bm{f},\bm{f}\rangle|\leq\varepsilon\|\bm{f}\|_{\bm{V}^{\prime}}^{2}\leq\varepsilon C\langle\mathcal{A}^{-1}\bm{f},\bm{f}\rangle,$
It follows that
$(1-\varepsilon C)\langle\mathcal{A}^{-1}\bm{f},\bm{f}\rangle\leq\langle\mathcal{M}_{\theta}\bm{f},\bm{f}\rangle\leq(1+\varepsilon C)\langle\mathcal{A}^{-1}\bm{f},\bm{f}\rangle.$
For sufficiently small $\varepsilon$, the coefficients $c_{1}$ and $c_{2}$ are positive and bounded, implying that the condition number of the preconditioned system is independent of the mesh. ∎
While theory suggests that neural operators can effectively approximate $\mathcal{A}^{-1}$, practical implementation encounters two significant limitations:
  • Spectral Bias: Empirical and theoretical studies indicate that neural networks trained via gradient descent learn low-frequency components significantly faster than high-frequency ones [rahaman2019spectral, xu2019frequency]. Consequently, when the target inverse operator contains rich high-frequency information, standard networks struggle to capture it efficiently.
  • Generalization Gap: The approximation accuracy is constrained by model capacity. Moreover, without architectures specifically designed for mesh invariance, performance often degrades when generalizing to unseen meshes or different refinement levels.
@@ 3.3 Hybrid Iterative Methods and Fourier Neural Solver
The analysis above highlights a complementarity: traditional smoothers (e.g., Weighted Block Jacobi) efficiently eliminate high-frequency errors but stagnate on low frequencies, whereas deep neural networks excel at capturing global low-frequency modes due to spectral bias. Leveraging this, we propose the DL-HIM
Smoothing Step: $\bm{u}^{(k+\frac{1}{2})}=\bm{u}^{(k)}+\mathcal{B}\bigl(\bm{f}-\bm{A}\bm{u}^{(k)}\bigr),$ (20a)
Correction Step: $\bm{u}^{(k+1)}=\bm{u}^{(k+\frac{1}{2})}+\mathcal{H}\bigl(\bm{f}-\bm{A}\bm{u}^{(k+\frac{1}{2})}\bigr).$ (20b)
Here, $\mathcal{B}$ denotes the traditional smoother, and $\mathcal{H}$ represents the neural network-based global corrector.
FNS [cui2022fourier] gives a specific realization of $\mathcal{H}$, inspired by the eigendecomposition $\bm{A}^{-1}=\bm{Q}^{-1}\Lambda^{-1}\bm{Q}$. For constant-coefficient PDEs on structured grids, the eigenvectors $\bm{Q}$ align with the Fourier basis, leading to the approximation
@@
where $\mathcal{F}$ is the fast Fourier transform and $\widetilde{\Lambda}$ is a learnable diagonal matrix. However, for general variable-coefficient problems, the true eigenbasis deviates from the Fourier basis. To address this, [cui2025hybrid] introduces a transition operator $\mathcal{C}$ to approximate the basis transformation $\bm{Q}\approx\mathcal{C}\mathcal{F}$
$\mathcal{H}\approx\bm{Q}^{*}\Lambda^{-1}\bm{Q}\approx\mathcal{F}^{-1}\,\mathcal{C}^{*}\,\widetilde{\Lambda}\,\mathcal{C}\,\mathcal{F}.$ (22)
Here, $\mathcal{C}$ is typically implemented as a convolutional neural network.
Despite improvements, FNS variants remain limited to structured grids and scalar PDEs due to their reliance on convolution operations. This work aims to generalize the FNS framework to support arbitrary unstructured meshes and vector-valued PDEs, overcoming previous geometric and dimensional constraints.
3.4 Convergence Analysis of Hybrid Iterative Methods
To demonstrate the theoretical robustness of the proposed method, we analyze the error propagation of the hybrid iteration involving the weighted Block Jacobi smoother $\mathcal{B}$ and the neural operator $\mathcal{H}$. The error propagation operator for one complete cycle is
$\mathcal{E}=(I-\mathcal{H}\mathcal{A})(I-\mathcal{B}\mathcal{A}).$ (23)
Our analysis employs the energy norm $\|\cdot\|_{a}$. As established by Korn’s inequality, this norm is equivalent to the $H^{1}$ norm, ensuring that the derived convergence rates are mesh-independent. We introduce an orthogonal decomposition of the space with respect to the energy inner product
$\bm{V}=\Theta^{\mathcal{B}}\oplus_{\perp_{a}}\Theta^{\mathcal{H}},$ (24)
where $\Theta^{\mathcal{B}}$ is the subspace of high-frequency errors effectively attenuated by the smoother, and $\Theta^{\mathcal{H}}$ is the low-frequency subspace targeted by the neural operator. For any error $\bm{e}\in\bm{V}$, we have the decomposition $\bm{e}=\bm{e}_{\mathcal{B}}+\bm{e}_{\mathcal{H}}$, satisfying
$\|\bm{e}\|_{a}^{2}=\|\bm{e}_{\mathcal{B}}\|_{a}^{2}+\|\bm{e}_{\mathcal{H}}\|_{a}^{2}.$ (25)
We characterize the operators with the following definitions and assumptions.
Definition 1 (Properties of the smoother).
The weighted Block Jacobi smoother $\mathcal{B}$ is considered effective if it acts contractively on $\Theta^{\mathcal{B}}$ while remaining stable on $\Theta^{\mathcal{H}}$. Specifically, there exists a constant $\rho\in[0,1)$ such that
$\|(I-\mathcal{B}\mathcal{A})\bm{v}\|_{a}$ $\leq\rho\,\|\bm{v}\|_{a},\,$ $\forall\,\bm{v}\in\Theta^{\mathcal{B}},$ (26)
$\|(I-\mathcal{B}\mathcal{A})\bm{v}\|_{a}$ $\leq\|\bm{v}\|_{a},\,$ $\forall\,\bm{v}\in\Theta^{\mathcal{H}}.$ (27)
Assumption 3.2 (Properties of the neural operator).
For the neural operator $\mathcal{H}$, we assume:
  1. 1.
    Low-Frequency Accuracy: For $\bm{v}\in\Theta^{\mathcal{H}}$, $\mathcal{H}$ approximates $\mathcal{A}^{-1}$ with high precision $\delta\ll 1$
    $\|(I-\mathcal{H}\mathcal{A})\bm{v}\|_{a}\leq\delta\|\bm{v}\|_{a}.$ (28)
  2. 2.
    High-Frequency Boundedness: For $\bm{v}\in\Theta^{\mathcal{B}}$, $\mathcal{H}$ is bounded by a constant $C\geq 1$
    $\|(I-\mathcal{H}\mathcal{A})\bm{v}\|_{a}\leq C\|\bm{v}\|_{a}.$ (29)
    The constant $C$ accounts for potential high-frequency amplification or noise introduction by the network.
Theorem 3 (Global convergence).
Let $\bm{e}^{(0)}$ be the initial error. The error after one hybrid iteration, $\bm{e}^{(1)}=\mathcal{E}\bm{e}^{(0)}$, satisfies
$\|\bm{e}^{(1)}\|_{a}\leq\gamma\|\bm{e}^{(0)}\|_{a},\qquad\text{where }\gamma=\sqrt{2}\max(\delta,\,C\rho).$ (30)
The hybrid iteration converges unconditionally provided the smoothing-boundedness condition $C\rho<\frac{\sqrt{2}}{2}$ holds.
Proof.
Decompose the initial error as $\bm{e}^{(0)}=\bm{e}^{(0)}_{\mathcal{B}}+\bm{e}^{(0)}_{\mathcal{H}}$. By (25), $\|\bm{e}^{(0)}\|_{a}=\sqrt{\|\bm{e}^{(0)}_{\mathcal{B}}\|_{a}^{2}+\|\bm{e}^{(0)}_{\mathcal{H}}\|_{a}^{2}}.$ The following proof is divided into three steps.
S1.
Smoothing Process
Applying the smoother (20a) yields the intermediate error
$\bm{e}^{(1/2)}=(I-\mathcal{B}\mathcal{A})\bm{e}^{(0)}.$
By linearity, we split this into $\bm{r}_{\mathcal{B}}=(I-\mathcal{B}\mathcal{A})\bm{e}^{(0)}_{\mathcal{B}}$ and $\bm{r}_{\mathcal{H}}=(I-\mathcal{B}\mathcal{A})\bm{e}^{(0)}_{\mathcal{H}}$. From Definition 1, we have
$\|\bm{r}_{\mathcal{B}}\|_{a}\leq\rho\|\bm{e}^{(0)}_{\mathcal{B}}\|_{a},\qquad\|\bm{r}_{\mathcal{H}}\|_{a}\leq\|\bm{e}^{(0)}_{\mathcal{H}}\|_{a}.$ (31)
Note: The operator action may disrupt orthogonality, so we treat $\bm{r}_{\mathcal{B}}$ and $\bm{r}_{\mathcal{H}}$ generally in the next step.
S2.
Correction Process
Applying the neural operator (20b) results in $\bm{e}^{(1)}=(I-\mathcal{H}\mathcal{A})\bm{e}^{(1/2)}$. Using the Triangle Inequality:
$\|\bm{e}^{(1)}\|_{a}$ $=\|(I-\mathcal{H}\mathcal{A})\bm{r}_{\mathcal{B}}+(I-\mathcal{H}\mathcal{A})\bm{r}_{\mathcal{H}}\|_{a}$
$\leq\|(I-\mathcal{H}\mathcal{A})\bm{r}_{\mathcal{B}}\|_{a}+\|(I-\mathcal{H}\mathcal{A})\bm{r}_{\mathcal{H}}\|_{a}.$
Invoking Assumption 3.2 (boundedness and accuracy) and the bounds from (31):
$\|\bm{e}^{(1)}\|_{a}$ $\leq C\|\bm{r}_{\mathcal{B}}\|_{a}+\delta\|\bm{r}_{\mathcal{H}}\|_{a}$
$\leq C\rho\|\bm{e}^{(0)}_{\mathcal{B}}\|_{a}+\delta\|\bm{e}^{(0)}_{\mathcal{H}}\|_{a}.$
S3.
Global Error Reassembly
Let $\gamma^{\prime}=\max(\delta,C\rho)$. Then $\|\bm{e}^{(1)}\|_{a}\leq\gamma^{\prime}(\|\bm{e}^{(0)}_{\mathcal{B}}\|_{a}+\|\bm{e}^{(0)}_{\mathcal{H}}\|_{a})$. Using the Cauchy-Schwarz inequality $(x+y)\leq\sqrt{2}\sqrt{x^{2}+y^{2}}$ for $x,y\geq 0$, combined with the initial orthogonality
$\|\bm{e}^{(1)}\|_{a}$ $\leq\gamma^{\prime}\sqrt{2}\sqrt{\|\bm{e}^{(0)}_{\mathcal{B}}\|_{a}^{2}+\|\bm{e}^{(0)}_{\mathcal{H}}\|_{a}^{2}}$
$=\sqrt{2}\max(\delta,C\rho)\|\bm{e}^{(0)}\|_{a}.$
Remark 1 (Energy Norm and Mesh Independence).
The choice of the energy norm $\|\cdot\|_{a}$ is physically significant. If the $L^{2}$ norm were used, the smoothing factor $\rho$ would degrade to $1$ and the amplification factor $C$ could tend to infinity as the mesh refines, due to the unboundedness of the differential operator $\mathcal{A}$. In contrast, Korn’s inequality ensures that $\rho$ and $C$ remain bounded mesh-independent constants under the energy norm, providing theoretical robustness for the multigrid-like behavior.
Remark 2 (Mode Mixing).
In problems with variable coefficients or irregular meshes, operators typically induce coupling between subspaces. The derivation explicitly accommodates this non-orthogonality by employing the triangle inequality during the intermediate steps. The result confirms that convergence is maintained as long as the smoother’s attenuation of high frequencies ($\rho$) sufficiently counteracts potential noise introduced by the neural operator ($C$).